Mickley Company’s predetermined overhead rate is $22.00 per direct labor-hour and its direct labor wage rate is $13.00 per hour. The following information pertains to Job A-500:

  

    
Direct materials $290
  Direct labor $195

  

Required:
1.

What is the total manufacturing cost assigned to Job A-500?

     

   

2.

If Job A-500 consists of 60 units, what is the average cost assigned to each unit included in the job? (Round your answer to 2 decimal places.)
Answer #1

1.
Direct Labor hour for Job A-500 = $195 / 13 = 15 DLHs

Overhead applied = 15 x $22 = $330

Total manufacturing cost assigned to Job A-500 = $290+195+330 = $815

2.
Average cost per unit = $815 / 60 = $13.58 per unit

If you have any query, kindly comment with your query and please mark thumbs up.